OpenStudy (anonymous):

(4/3)^-3 I'm supposed to rewrite it without an exponent, how?

OpenStudy (amistre64):

\[(\frac{a}{b})^{-n}=(\frac{b}{a})^{n}=\frac{b^n}{a^n}\]